OpenStudy (anonymous):

Find the integral of 1/x(x^2+4).....any ideas?

OpenStudy (dumbcow):

split it up using partial fractions \[\frac{1}{x(x^2+4)} = \frac{A}{x} + \frac{Bx}{x^2+4}\] \[\rightarrow A(x^2 +4) +Bx^2 = 1\] A+B = 0 4A = 1 ----> A = 1/4 ---> B = -1/4 Now the integral becomes \[\frac{1}{4}\int\limits \frac{dx}{x} - \frac{1}{4} \int\limits \frac{x}{x^2 +4} dx\]

OpenStudy (anonymous):

Thanks, I appreciate the start.

OpenStudy (dumbcow):

no problem